# Nagaoka ferromagnetism in semiconductor artificial graphene

### v2

@@ Abstract
We present the emergence of Nagaoka ferromagnetism in semiconductor-based artificial graphene with realistic Coulomb interaction using high-precision variational and diffusion Monte Carlo methods, complemented by exact diagonalization calculations of the extended Hubbard model. Specifically, we analyze a model of an armchair hexagonal geometry comprising $42$ lattice sites, nanopatterned on GaAs quantum wells with nearest-neighbor distance of $a=50$ nm. Our results reveal a distinct magnetic phase transition driven by the absence/addition of a single electron at half-filling where the ferromagnetic phase is further stabilized by Coulomb scattering terms. We determine effective Hubbard model parameters and identify the magnetic phase transition near $U/t\approx 60$.
@@
In this work, we predict the emergence of Nagaoka ferromagnetism in semiconductor AG, using advanced computational techniques. We show that Nagaoka ferromagnetism can be expected in systems with realistic Coulomb interactions, going beyond Hubbard model. Our findings not only deepen the fundamental understanding of strongly correlated electrons in synthetic lattices but also establish a framework for engineering novel quantum magnetic phases in artificial quantum materials. We employ continuum variational Monte Carlo (VMC) and diffusion Monte Carlo (DMC) methods within the fixed-node approximation to provide a non-perturbative and highly accurate treatment of many-body correlations [45, 46, 47, 48]. We focus on a hexagonal armchair geometry, which serves as an intermediary between finite-size samples and bulk graphene [44]. The system is modeled with a nearest-neighbor distance of $a=50$ nm, in alignment with recent experimental findings [27] and computational studies [44]. Our results reveal a transition from an antiferromagnetic (AFM) to a ferromagnetic (FM) phase which occurs exactly one hole/electron away from the half-filling as described in Nagaoka theorem [1]. Notably, this phase transition appears sharper when considering only the extremal values of $S_{z}$. However, a comprehensive scan over all $S_{z}$ values for a given potential depth and radius demonstrates a more gradual transition.
Our predictions are supported by appropriate exact diagonalization (ED) calculations of the Hubbard model on a hexagonal armchair lattice. Given the prohibitive computational complexity associated with the full Hilbert space for one hole added to the half-filling, the ED calculations are restricted to the subspace corresponding to a single spin flip from a fully spin polarized state. The resulting energy gains from this approach further support the phase transition identified through quantum Monte Carlo (QMC) simulations.
@@
For a deep understanding of transition dynamics including magnetic and metallic phases as a function of potential well depth (which controls $U/t$) and electron number, we consider a spin-spin correlation function averaged over all nearest neighbor pairs $(i,j)$, defined as $g=\langle m_{i}m_{j}\rangle/\langle n_{i}n_{j}\rangle$, where $m_{i}$ and $n_{i}$ are the average magnetization and electron density on site $i$ within a radius $r=a/2$. To reveal the internal spin structure, $m_{i}$ and $n_{i}$ are calculated using the pair densities $p_{\sigma\sigma_{0}}(\textbf{r},\textbf{r}_{0})$, the probability of finding an electron with spin $\sigma$ at location r when an electron with spin $\sigma_{0}$ is fixed at location $\textbf{r}_{0}$. The output values of the spin-spin correlation function remain in the $[-1,1]$ range, where $g=-1$ corresponds to an AFM phase, $g=0$ signifies a metallic configuration, and $g=1$ indicates a fully FM phase. In Fig. 1, spin-spin correlation function $g$ is shown as a function of potential depth $V_{0}$ for $N=42$ sites, $\rho=25$ nm and $k\neq 0$. The location of the fixed up-spin electron is denoted by a cross positioned atop a lattice site-strategically chosen to break system symmetry while avoiding edge effects. Both $N_{e}=42$ and $N_{e}=41$ electron numbers are considered for half-filling and a single hole away from half-filling, respectively (see supplemental materials (SM) for QMC $114$ sites results [49]). For these calculations, only maximum and minimum values of $S_{z}$ are considered to determine the ground state of the system. As shown in Fig. 1, the system at the half-filling exhibits a transition from metallic to AFM insulating phase around $V_{0}\approx-18$ meV, and the system with $N_{e}=41$ electrons shows a transition from metallic to FM phase around $V_{0}\approx-30$ meV. The results suggest that the system should make a transition to AFM phase before subtracting a single electron. If the same figure is analyzed vertically, the system exhibits a transition from AFM to FM phase around $V_{0}\approx-30$ meV when an electron is pulled away from the system as it is suggested in Nagaoka theorem. After $V_{0}\approx-30$ meV, the transition maintains its stability. Although, at $\rho=25$ nm, AG exhibits metallic-like behavior as predicted in [44], increasing effect of potential depth $V_{0}$ opens a way to have a localized behavior for the systems. To visualize the phases depicted above by $g$, Figs. 1a-b-c-d show the pair spin densities $p_{\uparrow\uparrow}(\textbf{r},\textbf{r}_{0})-p_{\downarrow\uparrow}(\textbf{r},\textbf{r}_{0})$. In Fig. 1(a), the system have a fully FM phase for $N_{e}=41$ electrons at $V_{0}=-45$ meV with maximum $S_{z}=41/2$ while the system with $N_{e}=42$ electrons at the same $V_{0}$ value Fig. 1(b) exhibits an AFM phase at minimum $S_{z}=0$. Additionally, at the lower extrema value of $V_{0}\approx-8.84$ meV, Fig. 1(c) and 1(d) shows metallic phases for $N_{e}=41$ with $S_{z}=1/2$ and $N_{e}=42$ with $S_{z}=0$, respectively.
@@
In summary, we demonstrated that Nagaoka ferromagnetism can be probed in semiconductor artificial graphene using accurate variational and diffusion Monte Carlo calculations, alongside exact diagonalization of the Hubbard model. Specifically, quantum Monte Carlo calculations on $N=42$ sites near half-filling reveal an antiferromagnetic to ferromagnetic phase transition driven by the absence of a single electron. Additionally, the phase transition can be induced by adding an electron to the half-filling, consistent with Nagaoka’s theorem on bipartite lattices. As $V_{0}$ decreases (i.e. $U/t$ increases), the $N_{e}=41$ electron system evolves from a liquid-like state to a fully ferromagnetic phase, passing through an intermediate antiferromagnetic phase where spin polarization gradually builds near the center of system, suggesting possible Nagaoka polaron formation. We estimate the $U/t$ ratio and determine that the phase transition occurs above $U/t\approx 60$, indicating that strong electron correlation is necessary. Furthermore, exact diagonalization of the Hubbard model, constrained to near-maximal polarization due to computational limitations, confirms the phase transition. Adding assisted-hopping and exchange interaction terms to the Hamiltonian stabilizes the transition that validates the $U/t$ ratio predictions of quantum Monte Carlo calculations. Thus, our findings establish the presence of an antiferromagnetic to ferromagnetic phase transition triggered by a single electron removal/addition in a solid-state system within a realistic Hamiltonian model fully accounting for long-range and exchange interactions.
@@ Acknowledgements.
The numerical calculations reported in this study were partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).

### v3

@@ Abstract
We present the emergence of Nagaoka ferromagnetism in semiconductor-based artificial graphene with realistic Coulomb interaction using high-precision variational and diffusion Monte Carlo methods, complemented by exact diagonalization calculations of the generalized Hubbard model. We analyze models of armchair hexagonal geometries nanopatterned on GaAs quantum wells. Our results reveal a distinct magnetic phase transition driven by the absence/addition of a single electron at half-filling. This form of itinerant magnetism predicted rigorously for Hubbard model remained unascertained in large scale realistic systems. We demonstrate that Coulomb scattering terms play a crucial role in stabilizing Nagaoka ferromagnetism, enabling the observation of the phase transition for system parameters near $U/t\approx 60$.
@@
In this work, we predict the emergence of Nagaoka ferromagnetism in semiconductor AG, using advanced computational techniques going beyond Hubbard model. We examine a 42-site hexagonal lattice — significantly larger than those studied in prior literature on Nagaoka ferromagnetism, and electrons interacting via long range Coulomb potential, using continuum variational Monte Carlo (VMC) and diffusion Monte Carlo (DMC) methods to provide a non-perturbative and highly accurate treatment of many-body correlations [45, 46, 47, 48]. We focus on a hexagonal armchair geometry, which serves as an intermediary between finite-size samples and bulk graphene, where edge effects are mitigated through a confining potential [44]. The system is modeled with a nearest-neighbor distance of $a=50$ nm, in alignment with recent experimental findings [27] and computational studies [44]. Our results reveal a transition from an antiferromagnetic (AFM) to a ferromagnetic (FM) phase which occurs exactly one hole/electron away from the half-filling as described in Nagaoka theorem [1]. Our predictions are supported by appropriate exact diagonalization (ED) calculations of the Hubbard model on a hexagonal armchair lattice. Given the prohibitive computational complexity associated with the full Hilbert space for one hole added to the half-filling, the ED calculations are restricted to the subspace corresponding to a single spin flip from a fully spin polarized state. The ED results further suggest that Coulomb interaction scattering terms are necessary to observe Nagaoka ferromagnetism for realistic system parameters.
@@
For a deep understanding of transition dynamics including magnetic and metallic phases as a function of potential well depth (which controls $U/t$) and electron number, we consider a spin-spin correlation function averaged over all nearest neighbor pairs $(i,j)$, defined as $g=\langle m_{i}m_{j}\rangle/\langle n_{i}n_{j}\rangle$, where $m_{i}$ and $n_{i}$ are the average magnetization and electron density on site $i$ within a radius $r=a/2$. To reveal the internal spin structure, $m_{i}$ and $n_{i}$ are calculated using the pair densities $p_{\sigma\sigma_{0}}(\textbf{r},\textbf{r}_{0})$, the probability of finding an electron with spin $\sigma$ at location r when an electron with spin $\sigma_{0}$ is fixed at location $\textbf{r}_{0}$. The output values of the spin-spin correlation function remain in the $[-1,1]$ range, where $g=-1$ corresponds to an AFM phase, $g=0$ signifies a metallic configuration, and $g=1$ indicates a fully FM phase. In Fig. 1, spin-spin correlation function $g$ is shown as a function of potential depth $V_{0}$ for $N=42$ sites, $\rho=25$ nm and $k\neq 0$. The location of the fixed up-spin electron is denoted by a cross positioned atop a lattice site-strategically chosen to break system symmetry while avoiding edge effects. Both $N_{e}=42$ and $N_{e}=41$ electron numbers are considered for half-filling and a single hole away from half-filling, respectively (see supplemental materials (SM) for QMC $N=114$ sites results [49]). For these calculations, only maximum and minimum values of $S_{z}$ are considered to determine the ground state of the system. As shown in Fig. 1, the system at the half-filling exhibits a transition from metallic to AFM insulating phase around $V_{0}\approx-18$ meV, and the system with $N_{e}=41$ electrons shows a transition from metallic to FM phase around $V_{0}\approx-30$ meV. The results suggest that the system should make a transition to AFM phase before subtracting a single electron. If the same figure is analyzed vertically, the system exhibits a transition from AFM to FM phase around $V_{0}\approx-30$ meV when an electron is pulled away from the system as it is suggested in Nagaoka theorem. After $V_{0}\approx-30$ meV, the transition maintains its stability. Although, at $\rho=25$ nm, AG exhibits metallic-like behavior as predicted in [44], increasing effect of potential depth $V_{0}$ opens a way to have a localized behavior for the systems. To visualize the phases depicted above by $g$, Figs. 1a-b-c-d show the pair spin densities $p_{\uparrow\uparrow}(\textbf{r},\textbf{r}_{0})-p_{\downarrow\uparrow}(\textbf{r},\textbf{r}_{0})$. In Fig. 1(a), the system have a fully FM phase for $N_{e}=41$ electrons at $V_{0}=-45$ meV with maximum $S_{z}=41/2$ while the system with $N_{e}=42$ electrons at the same $V_{0}$ value Fig. 1(b) exhibits an AFM phase at minimum $S_{z}=0$. Additionally, at the lower extrema value of $V_{0}\approx-8.84$ meV, Fig. 1(c) and 1(d) shows metallic phases for $N_{e}=41$ with $S_{z}=1/2$ and $N_{e}=42$ with $S_{z}=0$, respectively.
@@
In summary, we demonstrated that Nagaoka ferromagnetism can be probed in semiconductor artificial graphene using accurate variational and diffusion Monte Carlo calculations, alongside exact diagonalization of the Hubbard model. Specifically, quantum Monte Carlo calculations on $N=42$ sites near half-filling reveal an antiferromagnetic to ferromagnetic phase transition driven by the absence of a single electron. Additionally, the phase transition can be induced by adding an electron to the half-filling, consistent with Nagaoka’s theorem on bipartite lattices. As $V_{0}$ decreases (i.e. $U/t$ increases), the $N_{e}=41$ electron system evolves from a liquid-like state to a fully ferromagnetic phase, passing through an intermediate antiferromagnetic phase where spin polarization gradually builds near the center of system, suggesting possible Nagaoka polaron formation. We estimate the $U/t$ ratio and determine that the phase transition occurs above $U/t\approx 60$, indicating that strong electron correlation is necessary. Furthermore, exact diagonalization of the generalized Hubbard model, constrained to near-maximal polarization due to computational limitations, confirms the phase transition. Adding assisted-hopping and exchange interaction terms to the Hamiltonian stabilizes the transition that validates the $U/t$ ratio predictions of quantum Monte Carlo calculations. We note that while QMC calculations were performed for GaAs quantum dots parametrization, similar lattices can also be created in moiré superlattice [53]. Thus, our findings establish the presence of an antiferromagnetic to ferromagnetic phase transition triggered by a single electron removal/addition in a solid-state system within a realistic Hamiltonian model fully accounting for long-range and scattering interactions.
@@ Acknowledgements.
The quantum Monte Carlo calculations reported in this study were performed using the CHAMP program [54] and were partially carried out at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).